# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

### v1

@@ 3.1 Scheduler: Global and Local Time
We introduce the notion of time path in order to connect the local executions of robots in a distributed system to a global time variable. We consider that robots execute their protocols in a linearly ordered time set, either +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\;|\;x\geqslant 0\}$, or $\mathbb{N}$. Individually, each robot executes sequentially [^1] This is for simplicity and it should be possible to relax the assumption for a concurrent cycle at the cost of extra bookkeeping of how each step uses the information from the others to execute. the cycle of look, compute and move steps. We consider that cycles start with the move step, corresponding to an initial empty look and compute steps, so to match with the dynamical system definition.
The robots’ order of activation is defined by a scheduler, which chooses the robot that will execute the next step of its cycle. Common schedulers found in the literature are the synchronous, where all steps happens simultaneously, semi-synchronous, where the synchronization is guaranteed at every full cycle, and asynchronous, where there is no shared notion of time. Check [19] for further details on the taxonomy of schedulers. Any scheduler that is not fully synchronous may lead to global time that differs from the local clocks a robot may be equipped with. This distinction is made precise with the notion of a time path, matching the role of a global time.
[…]
@@
[…]
Furthermore, we suppose that $p$ is rectified in the sense that the length of the path [9] from $p(0)$ to $p(t)$ for all $t\in[0,T]$ has length $t$, using the infinity distance defined by $d_{\infty}((x_{1},\ldots,x_{n}),(y_{1},\ldots,y_{n}))=max(|x_{1}-y_{1}|,|x_{2}-y_{2}|,\ldots,|x_{n}-y_{n})$.
Given a time path $p$, the local time of robot $\mathbf{r}$ is given by $t_{\mathbf{r}}$ or $[t]_{\mathbf{r}}$, depending on if it is continuous or discrete.
Such a time path defines the relative progress of every robot at play, where a scheduler corresponds to a family of time paths. For instance, a synchronous scheduler can only correspond to a diagonal straight line, where all operations are aligned, while any path representing a semi-synchronous scheduler needs to cross all compute steps simultaneously.
Remark 3.2 (Notation for time).
It is important to note the different notations used for time. We use $t$ for the continuous time and $[t]$ for discrete time, as per Definition 2.1. A subscript with the robot identifier is used to depict local time, as in $t_{\mathbf{r}}$ or $[t]_{\mathbf{r}}$, and its absence represents global time.
We should note that, depending on the modeling choices, robots may or may not have access to a local clock that allows them to estimate or approximate the global time. Two possible executions of a multi-robot system composed of two robots is illustrated in Figure 1. In the case of instantaneous moves as in [1], because the behavior of robots would be undefined at points where a look operation (or event) would happen at the same time as a move operation, one should forbid points (represented as small squares in Figure 1) at the intersection of move and look events.
$[t]_{2}$$[t]_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{2}$$\textsc{l}_{2}$$\textsc{c}_{2}$$\textsc{m}_{3}$$\textsc{l}_{3}$$\textsc{c}_{3}$
@@
Figure 1: Two global time paths, blue and red, represent different possible executions of robots with local times $[t]_{1}$ and $[t]_{2}$. Each robot executes the sequence of actions (M) update ontic state, (L) observe environment and other robots, (C) update epistemic state. Each moment the time path crosses a dotted line, the associated robot executes the corresponding action. The actions are indexed by the cycle they belong to.
Note that communication scheduling is comprised in the current definition of the scheduler. Once the communication between robots is defined as part of the perception in the $\mathcal{LUMI}$ model or in Definition 2.5, a “delayed” message can be simplified to a robot not having executed the operation that would retrieve that information yet.
3.2 Executions of Distributed Multi-Robot Systems
Remarkably, the definitions of global and local times associated to a time path correspond precisely to the clock of the environment and the robot, respectively, in Definitions 2.9 and 2.10. This allows us to describe the robot executions, looking only at their local computations.
Definition 3.3 (Execution).
Let $\mathbf{r}\in\Pi$ be a robot, with an epistemic evolution specified by Definition 2.9. Let a pair $\langle\mathbf{e}_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}\rangle$ be a robot configuration. We say that an execution of $\mathbf{r}$ is a sequence $\{\langle\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\rangle\}$ indexed by its local time $[t]_{\mathbf{r}}\in N$, where $\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{r}}$, $\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{y}}$, and $\varphi_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}})=\mathbf{e}_{\mathbf{r}}[t+1]_{\mathbf{r}}$.
We can extend the notion of a robot execution to a system run, arriving at a global description of the system. We need first to define global robot configurations, which will take the role of possible worlds in the epistemic frame later described in Section 3.4.
Definition 3.4 (Global robot configuration).
A global robot configuration is tuple $\mathcal{C}=\langle\{\mathbf{e}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},\{\mathbf{y}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},[t]\rangle$ consisting of all local robot configurations at a global computation time $[t]$. $\mathcal{C}_{\mathbf{r}}$ is the restriction of a global configuration $\mathcal{C}$ to the robot $\mathbf{r}$.
Definition 3.5 (System run).
A system run $\sigma$ from time 00 to $T$ is a sequence of global robot configurations $\{\mathcal{C}[t]\}_{[t]\in[0,T]}$, where, for any $t\in[0,T]$:
  • there exists a subset $\varnothing\neq P(t)\subseteq\Pi$ of participating robots.
  • we have
    $\mathcal{C}_{\mathbf{r}}[t+1]=\\ \begin{cases}\varphi_{\mathbf{r}}(\mathcal{C}[t]_{\mathbf{r}})&\textrm{ if }\mathbf{r}\in P(t)\\ \mathcal{C}[t]_{\mathbf{r}}&\textrm{ if }\mathbf{r}\notin P(t)\end{cases}$
We denote the set of all possible runs $\mathcal{I}$, representing all the possible evolutions of the system solely based on the robot’s behavior. Note that this definition of run abstracts away the physical updates, in similar fashion to how the robot state machine represents the robots’ computing capacities. We effectively focus on the epistemic states, as the environment encompasses the interactions and constraints obtained from the ontic evolution and observations across all robots. The executions correspond to the individual epistemic evolution of a single robot, while the runs represent the epistemic evolution of all the robots at compatible cuts.
3.3 Solutions of Distributed Multi-Robot System
@@
We attest now the uniform treatment of time provided by the scheduler by bridging the understanding of system runs of robots as dynamical systems (Section 2.1) similarly to that of computing systems (Section 2.2), followed with a temporal-epistemic model pertinent to both perspectives in Section 4.
A distributed multi-robot system becomes an amalgamation of the individual robots obeying similar dynamics and epistemic update rules. The description of the global system is simplified by combining the state variables and observations as follows.
@@
Definition 3.6 (Distributed multi-robot system).
A distributed multi-robot system consists of $N$ robots, capable of observing the environment and interacting with the other robots through perception and communication. For a given time path $p$, its evolution is given by:
$\mathbf{U}[t]$ $=H(\mathbf{E}[t],t)$ (global control) (11)
This definition is a consequence of the evolution of single robots, as seen in Definitions 2.1 and 2.5. Changing variables in Equation 1, so that to consider the states of each robots at global time $t$ for some time path $p$, $\mathbf{X}(t)=\mathbf{x}(p_{i}(t))$, we get:
$\dot{\mathbf{X}}(t)$ $=\langle\dot{p},f(\mathbf{X}(t),\mathbf{u}[t])\rangle$
[…]
Equations (2), (3) and (4) now translate into, in terms of state $\mathbf{X}(t)$
$\mathbf{Y}(t)$ $=G(\mathbf{X}(t))$ (observation) (12)
$\mathbf{E}[t+1]$ $=\Phi(\mathbf{E}[t],\mathbf{Y}(t))$ (epistemic evolution) (13)
$\mathbf{U}[t]$ $=H(\mathbf{E}[t],t)$ (control) (14)
where $G=g$, $\Phi=\varphi$ and $H(\mathbf{E}[t],t)=h(\mathbf{E}[t],[t]_{\mathbf{r}})$.
@@
Consider the general differential inclusion
$\dot{x}\in\mathbb{F}(x)$ (15)
@@
In general, there can be many solutions to a differential inclusion. Throughout the section we note $S_{\mathbb{F}}(x_{0})$ the set of all (absolutely continuous) solutions to the Equation 15.
@@
Definition 3.8 (Marchaud map[2]).
The set-valued map $\mathbb{F}:\mathbb{R}^{n}\leadsto\mathbb{R}^{n}$ is a Marchaud map if $\mathbb{F}$ is upper semicontinuous with compact convex values and linear growth (that is, there is a constant $c>0$ such that $|\mathbb{F}(x)|:=sup\{|y|\ |\ y\in\mathbb{F}(x)\}\leqslant c(1+|x|)$, for every $x$).
We know from [2] that when $\mathbb{F}$ is a Marchaud map, then the inclusion (15) has a solution such that $x(t_{0})=x_{0}$ (for all $x_{0}$) and for a sufficiently small time interval $[t_{0},t_{0}+\varepsilon)$, $\varepsilon>0$. Global existence, for all $t\in\mbox{$\mathbb{R}$}$ can be shown provided $\mathbb{F}$ does not allow “blow-up” ($\|x(t)\|\rightarrow\infty$ as $t\rightarrow t^{*}$ for a finite $t^{*}$).
In that case, we can produce system runs from any solution $x$ of switched system (treated as a differential inclusion) of Equation (8) defined on time $[0,T]$, as in Definition 3.5 as follows. The system run is $\{\mathcal{C}(t)\}_{t\in[0,T]}$ is given by
$\mathcal{C}(t+1)_{\mathbf{r}}=\\ \begin{cases}G_{\mathbf{r}}(x(t),\eta(x(t)),\lambda(\mathbf{E}[t]))&\textrm{ if }\mathbf{r}\in P(t)\\ \mathcal{C}_{\mathbf{r}}(t)&\textrm{ if }\mathbf{r}\notin P(t)\end{cases}$
where $\mathbf{E}[t]$ is defined as in Equation (13).
3.4 System Runs and Epistemic Frames

### v2

@@ 3.1 Scheduler: Global and Local Time
The robots’ order of activation is given by a scheduler, which chooses the robot that will execute the next step of its cycle. Common schedulers found in the literature are the synchronous, where all steps happens simultaneously, semi-synchronous, where the synchronization is guaranteed at every full cycle, and asynchronous, where there is no shared notion of time. See [18] for more details on the taxonomy of schedulers. A scheduler that is not fully synchronous may lead to a global time that differs from the local time of the robots.
We introduce a time path to capture the relationship between the local execution of the robots and a common global time variable. A set of $n$ robots induces a space of possible executions $\mbox{$\mathbb{R}$}_{+}^{n}$, LCM cycles are mapped to each sequential unit interval. Any monotonic increasing path rooted in the origin of this space defines a possible execution order of the system. Each coordinate of $\mbox{$\mathbb{R}$}_{+}^{n}$ corresponds to the local time of a robot, and the length of the path corresponds to the global time. A robot is said to execute a step whenever a path’s projection to the corresponding local time axis crosses said step’s activation point. This leads to the following definition.
@@
For a system of $n$ robots, a time path is a differentiable map $p$, from $[0,T]$ to $\mbox{$\mathbb{R}$}^{n}_{+}$, for some $T\geq 0$, that has the following properties:
  • (initialisation) $p(0)=(0,\dots,0)$.
  • (monotonicity) For any $t,t^{\prime}\in[0,T]$ with $t\leq t^{\prime},p_{i}(t)\leq p_{i}(t^{\prime})$, for all components $i\in[1,\dots,n]$.
  • (rectification) For all $t\in[0,T]$, the global time $t$ is the maximum of all local times, i.e., $t=max(|p_{1}(t)-p_{1}(0)|,|p_{2}(t)-p_{2}(0)|,\dots,|p_{n}(t)\leavevmode\nobreak\ -\leavevmode\nobreak\ p_{n}(0)|)$.
A time path defines the relative progress of every robot, where the monotonicity property makes time always go forwards, and the rectification property sets the global time to match the fastest local time. A scheduler is then a family of time paths. For instance, a synchronous scheduler can only correspond to a diagonal straight line, where all operations are aligned. A semi-synchronous scheduler will have any of its possible time paths crossing all compute steps simultaneously, where synchronization happens. In the case of instantaneous moves, as in [1], the behavior of robots would be undefined at points where a look operation would happen at the same time as a move operation, and consequently one should forbid access to points near the intersection of move and look operations. Such forbidden zones (represented as small squares) and two possible executions of a multi-robot system composed of two robots is illustrated in Figure˜1.
$t]_{2}[$$t]_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{0}$$\textsc{l}_{0}$$\textsc{c}_{0}$$\textsc{m}_{1}$$\textsc{l}_{1}$$\textsc{c}_{1}$$\textsc{m}_{2}$$\textsc{l}_{2}$$\textsc{c}_{2}$$\textsc{m}_{3}$$\textsc{l}_{3}$$\textsc{c}_{3}$
@@
Figure 1: Two global time paths, blue and red, represent different possible executions of robots with local times $t]_{1}$ and $t]_{2}$. Each robot executes the sequence of actions (M) update ontic state, (L) observe environment and other robots, (C) update epistemic state. Each moment the time path crosses a dotted line, the associated robot executes the corresponding action. The squares are used to distinguish paths with different activation orders, as a look operation before and after a move provides different amounts of information. The actions are indexed by the cycle they belong to.
In the following, we use $t$ for the continuous time and $[t]$ for discrete time, as per Definition˜2.1. A subscript with the robot identifier is used to depict local time, as in $t_{\mathbf{r}}$ or $t]_{\mathbf{r}}$, and its absence represents global time. Given a time path $p$, the local time of robot $\mathbf{r}$ is given by $t_{\mathbf{r}}$ or $t]_{\mathbf{r}}$, depending on whether it is continuous or discrete.
3.2 Dynamical Runs of Distributed Multi-Robot System
@@
We attest now the uniform treatment of time provided by the scheduler by bridging the understanding of system runs of robots as dynamical systems (Section˜2.1) similarly to that of computing systems (Section˜2.2).
Given a time path $p$, the hybrid dynamics in global time coordinates $\mathbf{X}(t)=\mathbf{x}(p_{i}(t))$ for every robot $i$ satisfy
$\dot{\mathbf{X}}(t)=F_{p}(\mathbf{X}(t),\mathbf{U}[t])=\langle\dot{p},f(\mathbf{X}(t),\mathbf{U}[t])\rangle$
where $\langle\cdot,\cdot\rangle$ denotes the scalar product in $\mbox{$\mathbb{R}$}^{n}$. This connect the scheduler to the physical evolution via the time path.
@@
Consider the general differential inclusion $\dot{x}\in\mathbb{F}(x)$ where $\mathbb{F}$ is a map from $\mbox{$\mathbb{R}$}^{n}$ to $\mathcal{P}(R^{n})$, the set of subsets of $\mbox{$\mathbb{R}$}^{n}$. A function $x(\cdot):\mathbb{R}^{+}\rightarrow\mathbb{R}^{n}$ is a solution to the inclusion if $x$ is an absolutely continuous function and satisfies for almost all $t\in\mathbb{R}$, $\dot{x}(t)\in\mathbb{F}(x(t))$.
In general, there can be many solutions to a differential inclusion. Throughout the section we note $S_{\mathbb{F}}(x_{0})$ the set of all (absolutely continuous) solutions to the differential inclusion.
@@
We know from [2] that when $\mathbb{F}$ is a Marchaud map, then the inclusion has a solution such that $x(t_{0})=x_{0}$ (for all $x_{0}$) and for a sufficiently small time interval $t_{0},t_{0}+\varepsilon)$, $\varepsilon>0$. Global existence, for all $t\in\mbox{$\mathbb{R}$}$ can be shown provided $\mathbb{F}$ does not allow “blow-up” ($\|x(t)\|\rightarrow\infty$ as $t\rightarrow t^{*}$ for a finite $t^{*}$).
The solutions to the switched system in Definition˜2.3 provide the dynamical runs, treated as a differential inclusion defined on time $[0,T]$, as follows. The dynamical run is then the pair of trajectories $(\mathbf{X}(t),\mathbf{E}[t])$ for $t\in[0,T]$, representing the evolution of the global ontic and epistemic states generated by the switched system in Definition˜2.3.
3.3 Computational Runs of Distributed Multi-Robot Systems
We now describe the robot runs looking only at local epistemic computations. We use the fact that the global and local times associated to a time path correspond precisely to the clock of the environment and of the robots, respectively, in Definitions˜2.4 and 2.5.
Definition 3.3 (Execution).
Let $\mathbf{r}\in\Pi$ be a robot, with an epistemic evolution specified by Definition˜2.4. Let a pair $\langle\mathbf{e}_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}\rangle$ be a robot configuration. We say that an execution of $\mathbf{r}$ is a sequence $\{\langle\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\rangle\}$ indexed by its local time $t]_{\mathbf{r}}\in N$, where $\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{r}}$, $\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}}\in S_{\mathbf{y}}$, and $\varphi_{\mathbf{r}}(\mathbf{e}_{\mathbf{r}}[t]_{\mathbf{r}},\mathbf{y}_{\mathbf{r}}[t]_{\mathbf{r}})=\mathbf{e}_{\mathbf{r}}[t+1]_{\mathbf{r}}$.
The robot execution can now be extended to a computational system run, its robot system analog. For this, we define first the global robot configurations, which will take the role of possible worlds in the epistemic frame, later described in Section˜4.
Definition 3.4 (Global robot configuration).
A global robot configuration is a tuple $\mathcal{C}=\langle\{\mathbf{e}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},\{\mathbf{y}_{\mathbf{r}}\}_{\mathbf{r}\in\Pi},[t]\rangle$ consisting of all local robot configurations at a global computation time $[t]$. $\mathcal{C}_{\mathbf{r}}$ is the restriction of a global configuration $\mathcal{C}$ to the robot $\mathbf{r}$.
Definition 3.5 (System run).
A system run $\sigma$ from time 0 to $T$, determined by time path $p$, is a sequence of global robot configurations $\{\mathcal{C}[t]\}_{[t]\in[0,T]}$, where, for any $t\in[0,T]$:
  • there exists a subset $\varnothing\neq P_{p}(t)\subseteq\Pi$ of participating robots, activated by the time path $p$ at global time $t$.
  • we have
@@ 3.3 Computational Runs of Distributed Multi-Robot Systems
We denote the set of all possible runs $\mathcal{I}$, representing all the possible evolutions of the system solely based on the robot’s behavior. An execution correspond to the individual epistemic evolution of a single robot, while a run represent the epistemic evolution of all the robots. Note that this definition of run abstracts away the ontic (physical) updates in similar fashion to how the robot state machine represents solely the robots’ computations. We effectively focus on the epistemic states, while the ontic evolution related constraints and interactions are encompassed by the environment.
3.4 Correspondence of the Dynamical and Computational Perspectives
We establish now the conditions on which the computational model of a multi-robot system from Section˜2.2 is a faithful abstraction of the switched hybrid dynamical system of Section˜2.1. We show that the computational content of the system runs are the same in both dynamic (Section˜3.2) and computational (Section˜3.3) models. This justifies using the runs from computational models as the basis for the epistemic frames (Section˜4), as they capture all of the information available in the robots’ decision process.
We rely on a system abstraction procedure that converts a hybrid multi-robot system $\mathcal{H}$ into a computational system $\mathcal{C}$. We now define this abstraction and specify under which assumptions it preserves the computational information.
Definition 3.6 (System abstraction).
Given a hybrid system $\mathcal{H}$, a system abstraction $\Phi$ is a tuple of surjective functions $\Phi=(\Phi_{\mathbf{E}},\Phi_{\mathbf{Y}},\Phi_{\mathbf{U}},\Phi_{\mathbf{X}})$ that map the global state spaces as follows.
  • $\Phi_{\mathbf{E}}:\mathbf{E}\to S_{\mathbf{e}}$, partitions the global epistemic state space into a finite set of discrete memory states;
  • $\Phi_{\mathbf{Y}}:\mathbf{Y}\to S_{\mathbf{y}}$, maps the global observations into a finite set of discrete observations, as per Remark˜2.2;
  • $\Phi_{\mathbf{U}}:\mathbf{U}\to S_{\mathbf{u}}$, maps the global control commands to a finite set of global actions;
  • $\Phi_{\mathbf{X}}:\mathbf{X}\to S_{\mathcal{E}}$, partitions the global ontic state space into a finite set of environment states.
Lemma 3.7 (Induced run abstraction).
A system abstraction $\Phi:\mathcal{H}\to\mathcal{C}$ induces a function $\Phi^{*}:\Sigma_{\mathcal{H}}\to\Sigma_{\mathcal{C}}$ that maps hybrid runs to computational runs.
Proof 3.8.
The function $\Phi^{*}$ is constructed by applying the system abstraction to every trajectory $(\mathbf{X}(t),\mathbf{E}[t])\in\Sigma_{\mathcal{H}}$ at every discrete time step $t]\in[0,T]$.
The induced system abstraction need not be surjective, as some computational runs may not be physically possible. For instance, the computational model of a robot may allow for a state transition where its position updates faster than its speed limit. We filter those cases as follows.
Assumption 3.9 (Physical realizability).
Let $\mathcal{H}$ be a hybrid system with protocol defined by its epistemic evolution $\Phi$ and control function $CTL$. Let $\mathcal{C}$ be its computational abstraction. We assume that every run in $\Sigma_{\mathcal{C}}$ has a physical counterpart, i.e., $\Sigma_{\mathcal{C}}\subseteq\Phi^{*}(\Sigma_{\mathcal{H}})$.
Theorem 3.10 (Computational trace equivalence ).
Let $\Sigma_{\mathcal{H}}$ be the set of all runs of a computational system $\mathcal{H}$, and $\Sigma_{\mathcal{C}}$ the set of all runs of a hybrid system $\mathcal{C}$. If $\mathcal{C}$ is a faithful abstraction of a hybrid system $\mathcal{H}$, then $\Phi^{*}(\Sigma_{\mathcal{H}})=\Sigma_{\mathcal{C}}$.
Proof 3.11.
We proceed by showing the mutual inclusion $\Phi^{*}(\Sigma_{\mathcal{H}})\subseteq\Sigma_{\mathcal{C}}$ and $\Sigma_{\mathcal{C}}\subseteq\Phi^{*}(\Sigma_{\mathcal{H}})$.
$\Phi^{*}(\Sigma_{\mathcal{H}})\subseteq\Sigma_{\mathcal{C}}$. This follows from Lemma˜3.7, as the codomain of the induced system abstraction $\Phi^{*}$ is $\Sigma_{\mathcal{C}}$.
$\Sigma_{\mathcal{C}}\subseteq\Phi^{*}(\Sigma_{\mathcal{H}})$. This holds directly from assumption˜3.9.
This equivalence relies on assumption˜3.9, as we expect for the computation abstraction to be faithful to the physical dynamics. Theorem˜3.10 formalizes the consequence of this assumption.
4 Robot Tasks and Temporal Epistemic Logic